# On the role of the unitary transformations in Bell inequalities

### v1

@@ Abstract
The role of the unitary transformations in the Bell-CHSH inequality is highlighted, in both Quantum Mechanics and relativistic Quantum Field Theory. In the former case, we discuss, through a few examples, how the violation of the Bell-CHSH inequality can be encoded into unitary transformations acting on a given reference set of Bell’s observables. In the latter case, the Bell-CHSH inequality for a real massive scalar field in the vacuum state is considered. After introducing suitable bounded Hermitian operators, we show that using unitary transformations can significantly improve our previous numerical algorithm, leading to an increase in the size of the violation.
@@ I Introduction
The study of Bell inequalities within the contexts of Quantum Mechanics and Quantum Field Theory provides profound insights into the nature of quantum correlations and the interplay between locality, causality, and entanglement. In Quantum Mechanics, Bell inequalities serve as a fundamental test for the nonlocal correlations inherent in entangled states, providing a way to distinguish quantum predictions from those of classical theories rooted in local realism. Extensive theoretical investigations and experimental validations have consistently underscored the nonlocal nature of quantum mechanical systems [1, 2, 3, 4, 5, 6, 7, 8, 9, 10].
Extending the analysis of Bell inequalities to Quantum Field Theory introduces both significant challenges and new perspectives. The combination of Quantum Mechanics and Special Relativity creates a rich and sophisticated framework in which locality and causality form the foundational principles. Here, quantum fields are defined as operator-valued distributions. They interact with a vacuum state whose structure is deeply intricate, as illuminated by the Reeh-Schlieder theorem [11]. This complexity demands the use of advanced mathematical tools, including von Neumann Algebras, the Algebraic Quantum Field Theory and the Tomita-Takesaki modular theory, to rigorously examine entanglement and nonlocality in relativistic quantum systems [12, 13].
In the present work, we aim at pointing out the relevance of the unitary transformations in the study of the violation of a generalization of Bell inequalities, known as the Bell-Clauser-Horne-Shimony-Holt (Bell-CHSH) inequality [14, 1]. For a quantum bipartite system, the Bell-CHSH inequality reads:
@@
where $|\psi\rangle$ stands for the state of the system whose Hilbert space is ${\cal H}={\cal H}_{A}\otimes{\cal H}_{B}$. As it is customary, the letters $(A,B)$ refer to Alice and Bob.
A violation of the Bell-CHSH inequality occurs whenever
@@
$[A,B]=[A^{\prime},B]=[A,B^{\prime}]=[A^{\prime},B^{\prime}]=0\;,\qquad[A,A^{\prime}]\neq 0\;,\qquad[B,B^{\prime}]\neq 0\;.$ (4)
The possibility of making use of unitary transformations stems from the observation that the Bell operators are always determined up to unitary transformations, namely
@@
where $({\cal{U}}_{A},{\cal{U^{\prime}}}_{A})$ and $({\cal{U}}_{B},{\cal{U^{\prime}}}_{B})$ are unitaries acting on ${\cal H}_{A}$ and ${\cal H}_{B}$.
As one can easily figure out, these unitary operators might introduce new parameters, which can be adjusted so as to improve the size of the violation. As we shall see, this will be the case of the Bell-CHSH inequality for a relativistic scalar quantum field in the vacuum state.
The paper is organized as follows. In Section (II) we discuss the role of the unitaries in Quantum Mechanics by revisiting a few examples, namely: the spin $1/2$ Bell’s states, the $n00n$ states and the squeezed states. Section (III) is devoted to the Bell-CHSH inequality in Quantum Field Theory. The case of a real massive scalar field in $1+1$ Minkowski spacetime will be scrutinized. After introducing a set of bounded Hermitian operators, the violation of the Bell-CHSH inequality in the vacuum state is established by making use of the Tomita-Takesaki modular theory. Moreover, we show how the use of unitary transformation improves the size of the violation in our numerical setup. Section (IV) collects our conclusion.
@@ II Unitary transformations and the Bell-CHSH inequality in Quantum Mechanics
We begin this section by revisiting the textbook [^1] See [16] for an account on the Bell-CHSH inequality in Quantum Mechanics. example of the entangled spin $1/2$ singlet state
@@
As a second example, one might consider the $n00n$ states:
$|\psi_{n}\rangle=\frac{1}{\sqrt{2}}\left(|n_{A}\rangle\otimes|0\rangle_{B}+|0_{A}\rangle\otimes|n\rangle_{B}\right)\;,$ (16)
where $A,B$ correspond to the Hilbert spaces of two harmonic oscillators: ${\cal H}_{A}=\{|n_{A}\rangle=\frac{1}{\sqrt{n!}}({a^{\dagger}})^{n}|0_{A}\rangle\;,\quad n=0,1....\}$ and ${\cal H}_{B}=\{|n_{B}\rangle=\frac{1}{\sqrt{n!}}{({b^{\dagger}})^{n}}|0_{B}\rangle\;,\quad n=0,1....\}$ with
$[a,a^{\dagger}]=1\;,\qquad[b,b^{\dagger}]=1\;,\qquad[a,b]=[a,b^{\dagger}]=0\;.$ (17)
Here, the maximal violation of the Bell-CHSH inequality is achieved by taking as Bell’s observables the following expressions
$A$ $=$ $e^{i\alpha}|n_{A}\rangle\langle 0_{A}|\;+\;e^{-i\alpha}|0_{A}\rangle\langle n_{A}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{A}\rangle\langle m_{A}|\;,$
$A^{\prime}$ $=$ $e^{i\alpha^{\prime}}|n_{A}\rangle\langle 0_{A}|\;+\;e^{-i\alpha^{\prime}}|0_{A}\rangle\langle n_{A}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{A}\rangle\langle m_{A}|\;,$
$B$ $=$ $e^{i\beta}|n_{B}\rangle\langle 0_{B}|\;+\;e^{-i\beta}|0_{B}\rangle\langle n_{B}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{B}\rangle\langle m_{B}|\;,$
$B^{\prime}$ $=$ $e^{i\beta^{\prime}}|n_{B}\rangle\langle 0_{B}|\;+\;e^{-i\beta^{\prime}}|0_{B}\rangle\langle n_{B}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{B}\rangle\langle m_{B}|\;,$ (18)
where, as usual, $(\alpha,\alpha^{\prime},\beta,\beta^{\prime})$ denote the four Bell parameters. The maximal violation follows by noticing that
$\langle\psi_{n}|\;A\otimes B\;|\psi_{n}\rangle=\cos(\alpha+\beta)\;.$ (19)
Moreover, similarly to the case of spin $1/2$, the operators (18) can be cast into the form of Eq. (11), namely
$A={\cal U}^{\dagger}_{A}\;A_{0}\;{\cal U}_{A}\;,\qquad{\cal U}_{A}=e^{-i\alpha N_{A}}\;,\qquad N_{A}=a^{\dagger}a\;,$ (20)
while, for the reference operator, one gets
$A_{0}=|n_{A}\rangle\langle 0_{A}|\;+\;|0_{A}\rangle\langle n_{A}|\;+\;\sum_{m\neq 0,n}^{\infty}|m_{A}\rangle\langle m_{A}|\;.$ (21)
Analogous expressions can be written down for $(A^{\prime},B,B^{\prime})$. Again, all parameters $(\alpha,\alpha^{\prime},\beta,\beta^{\prime})$ are encoded into unitaries.
As a last example, let us consider the squeezed state
$|\lambda\rangle=(1-\lambda^{2})^{1/2}\sum_{n=0}^{\infty}\lambda^{n}|n_{A}n_{B}\rangle\;,\qquad|n_{A}n_{B}\rangle=|n_{A}\rangle\otimes|n_{B}\rangle\;,\qquad\lambda\in[0,1]\;.$ (22)
The Bell’s observables are introduced by making use of the isospin operators [16], i.e. by separating the basis states $\{|n_{A}\rangle,|n_{B}\rangle\}$ into even and odd states:
$|\lambda\rangle=(1-\lambda^{2})^{1/2}\sum_{n=0}^{\infty}\left(\;\lambda^{2n}|{(2n)}_{A}{(2n)}_{B}\rangle\;+\;\lambda^{(2n+1)}|{(2n+1)}_{A}{(2n+1)}_{B}\rangle\right)\;,$ (23)
and by defining
$A$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\alpha}|{(2n+1)}_{A}\rangle\langle(2n)_{A}|\;+\;e^{-i\alpha}|{(2n)}_{A}\rangle\langle(2n+1)_{A}|\right)\;,$
$A^{\prime}$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\alpha^{\prime}}|{(2n+1)}_{A}\rangle\langle(2n)_{A}|\;+\;e^{-i\alpha^{\prime}}|{(2n)}_{A}\rangle\langle(2n+1)_{A}|\right)\;,$
$B$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\beta}|{(2n+1)}_{B}\rangle\langle(2n)_{B}|\;+\;e^{-i\beta}|{(2n)}_{B}\rangle\langle(2n+1)_{B}|\right)\;,$
$B^{\prime}$ $=$ $\sum_{n=0}^{\infty}\left(\;e^{i\beta^{\prime}}|{(2n+1)}_{B}\rangle\langle(2n)_{B}|\;+\;e^{-i\beta^{\prime}}|{(2n)}_{B}\rangle\langle(2n+1)_{B}|\right)\;.$ (24)
It turns out that
$\langle\lambda|\;A\otimes B\;|\lambda\rangle=\frac{2\lambda}{1+\lambda^{2}}\cos(\alpha+\beta)\;,$ (25)
yielding maximal violation for $\lambda$ close to 1.
Again
$A$ $=$ ${\cal U}^{\dagger}_{A}\;A_{0}\;{\cal U}_{A}\;,\qquad{\cal U}_{A}=e^{-i\alpha N_{A}}\;,\qquad N_{A}=a^{\dagger}a\;,$
$A_{0}$ $=$ $\sum_{n=0}^{\infty}\left(\;|{(2n+1)}_{A}\rangle\langle(2n)_{A}|\;+\;|{(2n)}_{A}\rangle\langle(2n+1)_{A}|\;\right)\;.$ (26)
Other examples may be provided, including entangled coherent states [16] as well as Greenberger–Horne–Zeilinger (GHZ) states [16]. The above considerations show the relevant role that unitaries play in the violation of the Bell-CHSH inequality in Quantum Mechanics.
@@ III Facing the Bell-CHSH in Quantum Field Theory
We are now ready to face the more complex case of the relativistic Quantum Field Theory. We shall rely on the seminal works by S. J. Summers and R. Werner [12, 13, 17], who have established a set of remarkable results concerning the violation of the Bell-CHSH inequality in the vacuum state, see also the recent review [18].
III.1 The real massive scalar field in 1+1 Minkowski spacetime
Let us begin by considering a free massive real scalar field in $1+1$ Minkowski spacetime. In a plane-wave expansion, it reads
@@ III.1 The real massive scalar field in 1+1 Minkowski spacetime
where $\omega_{k}=k^{0}=\sqrt{k^{2}+m^{2}}$. For the canonical commutation relations, one has
[…]
@@
It is a well established fact that quantum fields have to be treated as operator-valued distributions [11]. As a consequence, they have to be smeared out to yield well-defined operators acting on the Hilbert space, i.e.
$\varphi(h)=\int\!d^{2}x\;\varphi(x)h(x),$ (29)
where $h$ is a real smooth test function with compact support. Using the smeared fields, the Lorentz-invariant inner product is introduced by means of the two-point smeared Wightman function
$\langle f|g\rangle$ $=\langle 0|\varphi(f)\varphi(g)|0\rangle=\frac{i}{2}\Delta_{PJ}(f,g)+H(f,g),$ (30)
where $f$ and $g$ are also real smooth test functions with compact support, and $\Delta_{PJ}(f,g)$ and $H(f,g)$ are the smeared versions of the Pauli-Jordan and Hadamard expressions
$\Delta_{PJ}(f,g)$ $=\int\!d^{2}xd^{2}yf(x)\Delta_{PJ}(x-y)g(y)\;,$
$H(f,g)$ $=\int\!d^{2}xd^{2}yf(x)H(x-y)g(y)\;.$ (31)
Here, $\Delta_{PJ}(x-y)$ and $H(x-y)$ are given by
$\Delta_{PJ}(t,x)$ $=$ $-\frac{1}{2}\;{\rm sign}(t)\;\theta\left(\lambda(t,x)\right)\;J_{0}\left(m\sqrt{\lambda(t,x)}\right)\;,$
$H(t,x)$ $=$ $-\frac{1}{2}\;\theta\left(\lambda(t,x)\right)\;Y_{0}\left(m\sqrt{\lambda(t,x)}\right)+\frac{1}{\pi}\;\theta\left(-\lambda(t,x)\right)\;K_{0}\left(m\sqrt{-\lambda(t,x)}\right)\;,$ (32)
where
$\lambda(t,x)=t^{2}-x^{2}\;,$ (33)
and $(J_{0},Y_{0},K_{0})$ are Bessel functions, while $m$ is the mass parameter.
Both the Hadamard and Pauli-Jordan distributions are Lorentz-invariant. Notably, the Pauli-Jordan distribution $\Delta_{PJ}(x)$ encodes relativistic causality, as it vanishes outside the light cone. Furthermore, $\Delta_{PJ}(x)$ and the Hadamard distribution $H(x)$ exhibit distinct symmetry properties: $\Delta_{PJ}(x)$ is odd under the transformation $x\rightarrow-x$, whereas $H(x)$ is even. When expressed in terms of smeared fields, the commutator of the field operators takes the form $\left[\varphi(f),\varphi(g)\right]=i\Delta_{PJ}(f,g)$. Within this context, the condition $\left[\phi(f),\phi(g)\right]=0,$ whenever the supports of $f$ and $g$ are spacelike separated, encapsulates causality in an elegant way.
@@ III.2 Basics of the Tomita-Takesaki modular theory
The Tomita-Takesaki modular theory serves as a powerful framework for analyzing the Bell-CHSH inequality in Quantum Field Theory [12, 13, 17]. To set the stage, it is worth briefly reviewing some of the basic features of this elegant theoretical structure.
@@
${\cal M^{\prime}}({\cal O})=\{g;\;\Delta_{PJ}(g,f)=0\;\;\forall f\in{\cal M}({\cal O})\}.$ (35)
@@
A very important property of the von Neumann algebra ${\cal W}({\cal M})$, Eq. (37), generated by the Weyl operators is that, due to the Reeh-Schlieder theorem [11, 19], the vacuum state $|0\rangle$ is both cyclic and separating, meaning that: $i)$ the set of states $\{{\cal A}_{h}|0\rangle$, ${\cal A}_{h}\in{\cal W}({\cal M})\}$ are dense in the Hilbert space; $ii)$ the condition $\{{\cal A}_{h}|0\rangle=0,{\cal A}_{h}\in{\cal W}({\cal M})\}$, implies ${\cal A}_{h}=0$.
@@
i.e., ${\cal W^{\prime}}({\cal M})$ contains all elements which commute with each element of ${\cal W}({\cal M})$. Let us also state the so-called Haag’s duality [19, 11, 20], ${\cal W^{\prime}}({\cal M})={\cal W}({\cal M^{\prime}}),$ namely, the commutant ${\cal W^{\prime}}({\cal M})$ coincides with the elements of ${\cal W}({\cal M^{\prime}})$ obtained by taking elements belonging to the symplectic complement ${\cal M^{\prime}}$ of ${\cal M}$. This duality, to our knowledge, has only been proven in the case of free Bose fields [20].
@@
We can now state the renowed Tomita-Takesaki theorem [21], namely:
@@
 $ii)$ there is a one-parameter family of operators $\Delta^{it}$ with $t\in\mathbb{R}$ which leave ${\cal W}({\cal M})$ invariant, that is:
@@
The operators $(s,s^{\dagger})$ have the physical meaning of projecting into the space ${\cal M}$ and its symplectic complement ${\cal M^{\prime}}$, namely, one can show [23, 13, 24] that a test function $f$ belongs to $\cal M$ if and only if
$sf=f\;.$ (52)
Analogously, a test function $g$ belongs to the symplectic complement ${\cal M^{\prime}}$ if and only if
@@
These properties enable us to construct a useful set of test functions for Alice and Bob. Following [12, 13], Alice’s test functions $(f,f^{\prime})$ can be specified as follows: Picking up the spectral subspace of $\delta$ specified by $[\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]\subset(0,1)$ and introducing the normalized vector $\phi$ belonging to this subspace, one writes
@@
it follows that the modular conjugation $j$ exchanges the spectral subspace $[\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]$ into $[1/\lambda^{2}-\varepsilon,1/\lambda^{2}+\varepsilon]$, ensuring that $\phi$ and $j\phi$ are orthogonal. Concerning now the pair of Bob’s test functions, they are given by $(jf,jf^{\prime})$. In fact, we have
@@
Finally, taking into account that $\phi$ belongs to the spectral subspace $[\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]$, it follows that [12, 13, 18], in the limit $\varepsilon\rightarrow 0$,
$||f||^{2}$ $=||jf||^{2}=\eta^{2}(1+\lambda^{2})$
$\langle f|jf\rangle$ $=2\eta^{2}\lambda$
$||f^{\prime}||^{2}$ $=||jf^{\prime}||^{2}={\eta^{\prime}}^{2}(1+\lambda^{2})$
$\langle f^{\prime}|jf^{\prime}\rangle$ $=2{\eta^{\prime}}^{2}\lambda$

### v2

# Bell Inequalities and Unitary Dynamics in Quantum Field Theory: Scalar and Proca Fields

@@ Abstract
We examine the role of unitary transformations in enhancing violations of the Bell-CHSH inequality in relativistic quantum field theory. In quantum systems, all relevant observables can be expressed via unitary rotations of reference operators. We extend this perspective to a real massive free scalar field and to the Proca vector field in $(1+1)$ Minkowski spacetime. In two dimensions, the Proca theory carries only one physical degree of freedom and is known to be equivalent to the scalar theory, implying that both yield identical Bell-CHSH violations. After introducing suitable bounded Hermitian operators, we employ Tomita–Takesaki modular theory to construct spacelike-separated observables in the vacuum. We demonstrate that unitary deformations significantly increase the Bell-CHSH violation, approaching the Tsirelson bound, and we report the first explicit evaluation of such violations in a vector-field vacuum.
@@ I Introduction
The study of Bell inequalities within the framework of quantum theory has been central to elucidating the profound distinction between quantum correlations and those permitted by local hidden variable theories. Originally formulated by Bell in 1964 [1], these inequalities place stringent constraints on statistical correlations predicted by any theory respecting locality and realism. Their violation, as predicted by quantum mechanics and experimentally confirmed with increasing sophistication [2, 3, 4, 5, 6, 7, 8, 9, 10, 11], provides compelling evidence for the intrinsically nonlocal character of entanglement.
While the formulation and experimental verification of Bell inequalities have traditionally been confined to non-relativistic quantum systems with finite degrees of freedom, their extension to the domain of relativistic quantum field theory (QFT) presents both conceptual and technical challenges. Unlike finite-dimensional Hilbert space settings, QFT inherently accommodates an infinite number of degrees of freedom and is governed by stringent demands of locality and covariance. Fields are operator-valued distributions, and their nontrivial vacuum structure (epitomized by the Reeh-Schlieder theorem [12]) precludes a naive transposition of quantum mechanical results into field-theoretic contexts.
To rigorously investigate entanglement and nonlocality in relativistic systems, it becomes necessary to employ the mathematical machinery of Algebraic Quantum Field Theory (AQFT), wherein local algebras of observables are associated with spacetime regions. Within this formalism, the Tomita–Takesaki modular theory plays a pivotal role, especially in light of Haag duality and the structure of von Neumann algebras generated by local field operators. These tools enable a principled construction of observables localized in spacelike separated regions, allowing for the definition of Bell-type inequalities directly in QFT. Crucial contributions by Summers and Werner [13, 14, 15] have shown that violations of the Bell-Clauser-Horne-Shimony-Holt (Bell-CHSH) inequality are not only possible in QFT but, remarkably, are generic even in the vacuum state.
For a non-relativistic quantum bipartite system, the Bell-CHSH inequality reads:
@@
where $|\psi\rangle$ stands for the state of the system whose Hilbert space is ${\cal H}={\cal H}_{A}\otimes{\cal H}_{B}$. As it is customary, the letters $(A,B)$ refer to Alice and Bob. A violation of the Bell-CHSH inequality occurs whenever
@@
$A,B]=[A^{\prime},B]=[A,B^{\prime}]=[A^{\prime},B^{\prime}]=0\;,\qquad[A,A^{\prime}]\neq 0\;,\qquad[B,B^{\prime}]\neq 0\;.$ (4)
In the present work, we revisit the Bell-CHSH inequality from the perspective of both quantum mechanics and relativistic QFT, with a particular focus on the role played by unitary transformations. In the quantum mechanical setting, we demonstrate through the representative example of the entangled spin-$1/2$ singlet states that the violation of the Bell-CHSH inequality can be encoded in unitary rotations acting on a fixed set of reference observables. The possibility of making use of unitary transformations stems from the observation that the Bell operators are always determined up to unitary transformations, namely
@@
where $({\cal{U}}_{A},{\cal{U^{\prime}}}_{A})$ and $({\cal{U}}_{B},{\cal{U^{\prime}}}_{B})$ are unitaries acting on ${\cal H}_{A}$ and ${\cal H}_{B}$. These unitary transformations introduce tunable parameters (often corresponding to physical angles or phases) that directly govern the extent of the violation.
Our central focus, however, lies in extending this analysis to relativistic QFT. We first consider the real massive scalar field in $(1+1)$ dimensions, employing the modular framework in Refs. [13, 14] to construct bounded Hermitian operators localized in wedge regions. Using these operators, we confirm the existence of a Bell-CHSH violation in the vacuum and show that unitary transformations substantially enhance its magnitude, pushing the correlator value closer to Tsirelson’s bound.
A novel contribution of this paper is the extension of the analysis to the Proca vector field in $(1+1)$ dimensions. In higher dimensions, a Proca field propagates massive spin-1 excitations with multiple polarization states. However, in two dimensions the transversality condition eliminates any independent polarization, leaving only a single propagating degree of freedom. More precisely, the Proca field is dually equivalent to a massive scalar field: one can represent the field strength as $F_{\mu\nu}=m^{2}\,\varepsilon_{\mu\nu}\,\varphi$, or, equivalently, the vector potential as $A^{\mu}=\varepsilon^{\mu\nu}\partial_{\nu}\varphi$, where $\varphi$ is a scalar field and $\varepsilon_{\mu\nu}$ is the two-dimensional Levi-Civita symbol. [^1] This duality is exact at the level of free field theories and ensures that the physical Hilbert spaces of the Proca and scalar models are isomorphic. The massive scalar field may thus be regarded as the unique dynamical degree of freedom encoded in the Proca field in 1+11+1D. By carefully implementing the smearing procedure with transverse test functions and exploiting this equivalence, we show that the Bell-CHSH violation in the Proca vacuum reproduces exactly the scalar case, both before and after the introduction of unitary transformations. To the best of our knowledge, this constitutes the first explicit computation of Bell-CHSH violations for a vector field in the QFT vacuum, and it highlights how dualities between field theories can directly shape the manifestation of quantum nonlocality.
The paper is organized as follows. In Section II, we review the role of unitary transformations in the Bell-CHSH inequality within quantum mechanics. Section III introduces the modular-theoretic construction of localized operators in QFT and applies it to the real scalar field. Section IV extends the analysis to the Proca field in $(1+1)$ dimensions. Our conclusions and perspectives for future developments are presented in Section V.
@@ II Unitary transformations and the Bell-CHSH inequality in Quantum Mechanics
We begin this section by revisiting the textbook [^2] See [17] for an account on the Bell-CHSH inequality in Quantum Mechanics. example of the entangled spin-$1/2$ singlet state
@@
Other examples may be provided, including entangled coherent states, Greenberger-Horne-Zeilinger (GHZ) states, the $N00N$ states and the squeezed state [17, 18]. The above considerations show the relevant role that unitaries play in the violation of the Bell-CHSH inequality in Quantum Mechanics.
@@ III Facing the Bell-CHSH in Quantum Field Theory
We are now ready to face the more complex case of the relativistic QFT. We shall rely on the original works by S. J. Summers and R. Werner [13, 14, 15], who have established a set of remarkable results concerning the violation of the Bell-CHSH inequality in the vacuum state, see also the recent review [19].
III.1 The real massive scalar field in $(1+1)$ Minkowski spacetime
We consider a free real scalar field of mass $m$ in $(1+1)$-dimensional Minkowski spacetime. Its quantized form, expressed via plane-wave decomposition, reads
@@ III.1 The real massive scalar field in $(1+1)$ Minkowski spacetime
where $\omega_{k}=\sqrt{k^{2}+m^{2}}$ and the annihilation and creation operators satisfy the canonical commutation relations
$[a_{k},a^{\dagger}_{q}]$ $=2\pi\,2\omega_{k}\,\delta(k-q),$ (17)
@@
In a rigorous QFT framework, the scalar field is treated as an operator-valued distribution [12], and thus must be smeared with smooth, compactly supported test functions to yield well-defined operators on the Hilbert space. Given a test function $h({\bf\it x})\in C_{0}^{\infty}(\mathbb{R}^{1,1})$, with ${\bf\it x}=(t,x)$, the corresponding smeared field operator is defined as
$\varphi(h)=\int\!d^{2}{\bf\it x}\;\varphi({\bf\it x})h({\bf\it x}).$ (18)
The Lorentz-invariant vacuum expectation value of two smeared field operators is defined through the two-point Wightman function:
$\langle f|g\rangle$ $=\langle 0|\varphi(f)\varphi(g)|0\rangle=\frac{i}{2}\Delta_{\rm PJ}(f,g)+H(f,g),$ (19)
where $f,g\in C_{0}^{\infty}$, and $\Delta_{\text{PJ}}(f,g)$ and $H(f,g)$ denote the smeared Pauli–Jordan and Hadamard distributions, respectively. These are defined by
$\Delta_{\rm PJ}(f,g)$ $=\int\!d^{2}{\bf\it x}d^{2}{\bf\it y}f({\bf\it x})\Delta_{\rm PJ}({\bf\it x}-{\bf\it y})g({\bf\it y})\;,$
$H(f,g)$ $=\int\!d^{2}{\bf\it x}d^{2}{\bf\it y}f({\bf\it x})H({\bf\it x}-{\bf\it y})g({\bf\it y})\;.$ (20)
with distributional kernels given explicitly by
$\Delta_{\rm PJ}({\bf\it x})$ $=$ $-\frac{1}{2}\;{\rm sign}(t)\;\theta\left(\lambda({\bf\it x})\right)\;J_{0}\left(m\sqrt{\lambda({\bf\it x})}\right)\;,$
$H({\bf\it x})$ $=$ $-\frac{1}{2}\;\theta\left(\lambda({\bf\it x})\right)\;Y_{0}\left(m\sqrt{\lambda({\bf\it x})}\right)+\frac{1}{\pi}\;\theta\left(-\lambda({\bf\it x})\right)\;K_{0}\left(m\sqrt{-\lambda({\bf\it x})}\right)\;,$ (21)
where the Lorentz-invariant quantity
$\lambda({\bf\it x})=t^{2}-x^{2}$ (22)
determines the causal structure. Here, ($J_{0}$, $Y_{0}$, $K_{0}$) are Bessel functions of the first kind, second kind, and modified second kind, respectively.
The Pauli-Jordan distribution $\Delta_{\rm PJ}({\bf\it x})$ is Lorentz-invariant and encodes causality, vanishing outside the light cone. Moreover, it is odd under ${\bf\it x}\rightarrow-{\bf\it x}$, whereas the Hadamard function $H({\bf\it x})$ is even. The commutator of the field operators is thus $\left[\varphi(f),\varphi(g)\right]=i\Delta_{\rm PJ}(f,g)$, ensuring that $\left[\phi(f),\phi(g)\right]=0,$ whenever the supports of $f$ and $g$ are spacelike separated. This compactly encodes the principle of microcausality in relativistic field theory.
@@ III.2 Basics of the Tomita-Takesaki modular theory
The Tomita-Takesaki modular theory serves as a powerful framework for analyzing the Bell-CHSH inequality in QFT [13, 14, 15]. To set the stage, it is worth briefly reviewing some of the basic features of this elegant theoretical structure.
@@
${\cal M^{\prime}}({\cal O})=\{g;\;\Delta_{\rm PJ}(g,f)=0\;\;\forall f\in{\cal M}({\cal O})\}.$ (24)
@@
A very important property of the von Neumann algebra ${\cal W}({\cal M})$, Eq. (26), generated by the Weyl operators is that, due to the Reeh-Schlieder theorem [12, 20], the vacuum state $|0\rangle$ is both cyclic and separating, meaning that: $i)$ the set of states $\{{\cal A}_{h}|0\rangle$, ${\cal A}_{h}\in{\cal W}({\cal M})\}$ are dense in the Hilbert space; ii)ii) the condition $\{{\cal A}_{h}|0\rangle=0,{\cal A}_{h}\in{\cal W}({\cal M})\}$, implies ${\cal A}_{h}=0$.
@@
[…]
@@
We can now state the renowned Tomita-Takesaki theorem [22], namely:
@@
 ii)ii) there is a one-parameter family of operators $\Delta^{it}$ with $t\in\mathbb{R}$ which leave ${\cal W}({\cal M})$ invariant, that is:
@@
[…]
sf=f.sf=f\;. (41)
[…]
@@
These properties enable us to construct a useful set of test functions for Alice and Bob. Following [13, 14], Alice’s test functions $(f,f^{\prime})$ can be specified as follows: Picking up the spectral subspace of $\delta$ specified by $\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]\subset(0,1)$ and introducing the normalized vector $\phi$ belonging to this subspace, one writes
@@
it follows that the modular conjugation $j$ exchanges the spectral subspace $\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]$ into $1/\lambda^{2}-\varepsilon,1/\lambda^{2}+\varepsilon]$, ensuring that $\phi$ and $j\phi$ are orthogonal. Concerning now the pair of Bob’s test functions, they are given by $(jf,jf^{\prime})$. In fact, we have
@@
Finally, taking into account that $\phi$ belongs to the spectral subspace $\lambda^{2}-\varepsilon,\lambda^{2}+\varepsilon]$, it follows that [13, 14, 19], in the limit $\varepsilon\rightarrow 0$,
[…]
$\langle f|jf\rangle$ $=2\eta^{2}\lambda,$
[…]
$\langle f^{\prime}|jf^{\prime}\rangle$ $=2{\eta^{\prime}}^{2}\lambda,$